OpenStudy (anonymous):

Fans & Medals!! Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A. x + y = y + 5 B. x = (x – 6) + 5 C. x + (y + 5) = 6 D. (y + 5) + y = 6

OpenStudy (dtan5457):

Which of these isolate x or y?

OpenStudy (dtan5457):

x + y = 6 x = y + 5 Which of these plug the 2nd one (isolation of x) into the 1st equation?

OpenStudy (anonymous):

?? Not good at this stuff explain plz @dtan5457

OpenStudy (dtan5457):

When you solve a system by substitution.... You take y=whatever or x=whatever to plug into the 2nd equation.

OpenStudy (dtan5457):

x + y = 6 x = y + 5 you know that x=y+5 plug that into x+y=6

OpenStudy (dtan5457):

and you get...?

OpenStudy (anonymous):

c?

OpenStudy (dtan5457):

your so close but no, it's not C. in C, they substituted for the value of Y, not x.

OpenStudy (anonymous):

Ok so D?